The purpose of the capacitor lab was to find C1 and C2 and to verify the equations for equivalent capacitance for capacitors in series and in parallel. For the capacitor in parallel, I got a relative error of 3%. For the capacitor in series, i got a relative error of 22%. 22% seems too large of a number for me to be able to verify the equation 1/C_eq= 1/C1 + 1/C2. Any thoughts on weather 22% is too large or not?
Answer #1

22 % is a very large error .

In parallel Ceq = C1 + C2   , and with 3% error let us say we get Ceq \pm 0.03 Ceq ,

In parallel the amount of error is carried by 2 capacitors is summed up in the Equivalent capacitence

but in series connection I / Ceq = 1 / C1 + 1 / C2

Ceq = ( C1 * C2) / ( C1 + C2) which implies the error even gets reduced accordingly .

So the error in series connection should be the minimum.